OpenStudy (anonymous):

FAN AND MEDAL AND FAN TESTIMONIAL TO WHOEVER HELPS!!!! Solve 64x = 16^(x−1) Idk how to do this please show me how?!

OpenStudy (anonymous):

@Abdulhameed @jdoe0001 @hartnn

OpenStudy (anonymous):

check your question.

OpenStudy (anonymous):

r u sure about the question ?

OpenStudy (anonymous):

how?! that's what it says.... \[64x = 16^{x - 1}\]

OpenStudy (anonymous):

yeah.....

OpenStudy (anonymous):

ok but what did u write in the previous comment bcoz it say mathh processing error

OpenStudy (anonymous):

so 64x = 16^(x-1) the whole x - 1 thing is an exponent

OpenStudy (anonymous):

if it was 16 not 16x it would be very easy , the problem is the x in 16x

OpenStudy (anonymous):

hmmm...

OpenStudy (anonymous):

i think may be it's 16^x not 16x

OpenStudy (anonymous):

i have this thing here

OpenStudy (anonymous):

the answer choice i picked was random.... don't bother with it

OpenStudy (anonymous):

oh it's 64^x=16^x-1

OpenStudy (anonymous):

wait.... i didn't put that?!?!?!

OpenStudy (anonymous):

oooopppsss nevermind, i didn't ... sorry

OpenStudy (anonymous):

\[64^x=16^{x-1}\] \[\left( 4^3 \right)^x=\left( 4^2 \right)^{x-1},4^{3x}=4^{2x-2}\] 3x=2x-2 x=?

OpenStudy (anonymous):

64=4^3 16 =4^2 64^x=4^3x 16^x-1=4^2x-2 4^3x=4^2x-2 3x=2x-2 3x-2x=-2 x=-2

OpenStudy (anonymous):

thanks to both of you! :D

OpenStudy (anonymous):

u r very welcome (^_^)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

hey @HelpBlahBlahBlah .... if ur online... can u give @surjithayer a medal? i can only give one person a medal..... but both helped me a lot thx :D

OpenStudy (helpblahblahblah):

yea sure

OpenStudy (anonymous):

kk thanks! :D

OpenStudy (helpblahblahblah):

anytime :)

OpenStudy (anonymous):

:D

OpenStudy (helpblahblahblah):

lol :P

OpenStudy (anonymous):

hey @Abdulhameed you still there?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

i wanna ask... what if the number on the first side is a fraction? like in...

OpenStudy (anonymous):

Solve 1/36 = 6^(x−3)

OpenStudy (anonymous):

u have to make the numbers that have the exponents equal 1/36=1/6^2=6^-2 6^-2=6^(x-3) -2=x-3 -2+3=x x=1

OpenStudy (anonymous):

ohhhh... ok Thanks again :D

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

@31356

OpenStudy (31356):

Correct! :D